# ARKit LabelMaker: A New Scale for Indoor 3D Scene Understanding

### v1

@@ 4.2 Datasets and Metrics for Evaluation
Table 1: The size of dataset that is used for training and evaluation in this work. We provide by far the largest real-world labeled training dataset compared to existing real-world datasets. We provide automatically generated dense semantic annotations for 4471 training trajectories and 548 validation trajectories.
4.3 Experiment Settings
Pre-training. To investigate whether automatic labels are useful to learn strong features from imperfect annotations, we pre-train both, Minkowskinet [7] and PointTransformerV3, on our generated ALS200 dataset. Afterwards, we fine-tune the pretrained models on the ScanNet and ScanNet200 dataset, respectively.
For PTv3, we adhere to the settings outlined in [36] employing the AdamW optimizer with OneCycleLR for 800 epochs of training. However, the learning rate during pre-training is adjusted to 0.0016. Similar to the fine-tuning of MinkowskiNet, we initially freeze the backbone during fine-tuning and solely train the classification head until convergence. Then, we fine-tune on ScanNet or ScanNet200 with a reduced learning rate of 0.0006. Besides ALS200, we also pre-train PTv3 with ALC label space as the mapping from wordnet to ScanNet200 may reduce neural stimulation.
Co-training with ALS200. With this experiment, we aim to investigate if our ALS200 dataset can be simply combined with existing datasets for increasing the dataset size and, therefore, the performance of the resulting model. To this end, we combine ALS200 with the ScanNet200 dataset and train a MinkwoskiNet from scratch. Due to resource limitations, we could only train MinkowskiNet with this combined dataset. The training setting is exactly same as the pre-training stage of MinkowskiNet described above.
Joint-training. We employ PTv3+PPT for joint training on multiple datasets across multiple label spaces. Besides ScanNet/ScanNet200, ScanNet++, S3DIS and Structured3D, we add our ALC dataset. We choose LabelMaker’s wordnet label space in order to provide the network with maximum possible semantic class stimulation. We use the exact same setting of PTv3+PPT from [36]. We use AdamW optimizer with OneCycleLR scheduler and a learning rate of 0.05. Additionally, we incorporate the LabelMaker WordNet label space into the norm layer and the final classification head.
4.4 Results
@@
fine-tune (Ours) ALS200 → ScanNet 77.0 -
@@
fine-tune (Ours) ALS200 → ScanNet 81.2 -
fine-tune (Ours) ALC → ScanNet 80.6 79.0
@@
PPT (Ours) ScanNet+ ScanNet200 + ScanNet++ + Structure3D + ALC 81.1 79.8
@@
Table 2: Semantic Segmentation Scores on ScanNet20. We compare different training strategies for two top-performing models (PointTransformerv3 [36] and MinkowskiNet [7]) on the ScanNet20 dataset. We can show for both models adding ALS200 through pre-training and co-training improves the performance for both models. With PonderV2 [42] and Mix3D [20], we compare large-scale pretraining to two other training strategies. We can show that large-scale pre-training is superior to both, extensive data augmentation (Mix3D) and self-supervised pre-training (PonderV2).
In Table 2, we present the results for the ScanNet dataset. In the case of MinkowskiNet [7], we can not only show that pre-training on our large-scale, real-world ALS200 dataset significantly improves the mean intersection-over-union compared to vanilla training, but it also significantly outperforms to other variants of pre-training. Pre-training on our imperfect yet automatically generated labels is superior to self-supervised pre-training (PonderV2 [42]) and extensive data augmentation (Mix3D [20]). This indicates that direct supervision with scale in training data for supervised learning is essential for 3D semantic segmentation. Additionally, the ALS200/ALC pre-trained PTv3 exhibits comparable or superior performance to large-scale multi-dataset joint training. This proves that our dataset is more data-efficient compared with synthetic dataset.
Method Training Data val test
MinkUNet [7]
vanilla ScanNet200 29.3 25.3
fine-tune (Ours) ALS200 → ScanNet200 30.1 27.4
co-training (Ours) ALS200 + ScanNet200 30.6 -
PTv3 [36]
vanilla ScanNet200 35.2 37.8
fine-tune (Ours) ALS200 → ScanNet200 38.4 -
fine-tune (Ours) ALC200 → ScanNet200 38.7 38.4
PPT [36] ScanNet200 + S3DIS + Structure3D → ScanNet200 36.0 39.3
PPT(Ours) ScanNet+ ScanNet200 + ScanNet++ + Structure3D + ALC 40.3 41.4
Table 3: Semantic Segmentation Scores on ScanNet200 [29]. To investigate our large-scale dataset also helps with long-tail categories, we evaluate it on the ScanNet200 dataset. For both, MinkowskiNet [7] and PointTransformerv3 [36], we compare it to vanilla training as well as training procedure proposed in [36]. We can show that common neural networks benefit from pre-training on automatically generated large-scale annotations compared.
Similar trends are also visible in Table 3, where we show the results on the ScanNet200 dataset. The co-training of MinkowskiNet on the ScanNet200 dataset demonstrates that our ALS200 dataset facilitates training without inducing a domain gap in comparison to the ScanNet200 dataset.
We also incorporate our ARKit LabelMaker dataset into the joint training for PTv3+PPT, leading to a significant boost in validation and test mIoU. This version of PTv3 achieves state-of-the-art performance on both ScanNet and ScanNet200, with the latter benefiting greatly from improved tail class mIoU( Table B1).Our dataset’s effectiveness serves as a good complementary for rare classes in ScanNet200.
In Table 4 we show the results on ScanNet++ dataset. As on the other datasets, pre-training on ALC improves validation set performance compared to vanilla PTv3 training on ScanNet++. However, test set performance deteriorates. For joint training, we conducted two trials: one with a small and one with a larger training scale. In the first, we excluded Structured3D, as PTv3’s updated version (28,212 samples) was too large for our machines. In the second trial, we use the old version of Structured3D which has 8216 samples and followed Wu et al.’s suggestion to exclude S3DIS due to its high memory consumption. These results suggest that the ALC dataset may have limited benefits for ScanNet++ training.
PTv3 Variant Training Data #Data val mIoU test top-1/3 mIoU
vanilla ScanNet++ 713 41.8 45.8 / 69.7
fine-tune (Ours) ALC200 → ScanNet++ 4471 → 713 42.5 43.7 / 65.5
PPT [36] ScanNet200 + ScanNet++ + Structure3D 45868 45.3 46.5 / 71.1
PPT (Ours) ScanNet200 + ScanNet++ + ALC 11168 44.5 46.1 / 70.8
PPT (Ours) ScanNet+ ScanNet200 + ScanNet++ + Structure3D + ALC 30386 44.6 46.1 / 68.5

### v2

@@ 4.2 Datasets and Evaluation Metrics
Table 1: Dataset Size. We provide by far the largest real-world labeled training dataset compared to existing real-world datasets. We provide automatically generated per-point semantic annotations for 4471 training scenes and 548 validation scenes.
ScanNet [9]
comprises 1513 densely annotated scans across 707 distinct indoor scenes, totaling 2.5 million RGB-D frames. It stands as one of the most widely used and influential benchmark datasets for indoor 3D scene understanding. It is annotated by humans using the NYU40 label space and evaluated on a subset of 20 classes from NYU40.
ScanNet200 [30].
While only 20 classes are used in the ScanNet benchmark, the original dataset is annotated with many more classes. ScanNet200 [30] leverages these annotations and organizes them into a new benchmark with 200 classes that are of higher-resolution than the original ScanNet classes. Given the large-number of different categories generated by our LabelMakerV2 pipeline, we also pre-train the models for this task and evaluate them on the ScanNet200 benchmark.
ScanNet++ [43]
is a dataset of 460 high-resolution 3D indoor scenes with dense semantic and instance annotations, captured using a high-precision laser scanner and registered images from a DSLR camera and RGB-D streams. It focuses on long-tail and multi-labeled annotations. Models are typically evaluated on 100 classes.
Structured3D [48]
is a large-scale indoor synthetic RGB-D dataset featuring 6519 training scenes and 1697 test scenes. It is annotated with a label space of 25 classes. Structured3D is only used in PTv3+PPT joint training and we adopt pre-processed version of these two datasets from [40].
Matterport3D [6]
is a large-scale RGB-D dataset containing 10,800 panoramic views from 194,400 RGB-D images across 90 building-scale scenes. We map the ScanNet200 label space to Matterport3D and use this dataset for zero-shot evaluation of our trained model.
ARKit LabelMaker (Ours).
This is the dataset generated with our method described above. The resulting dataset contains 5019 scenes, from which we take 4471 for training and 548 for validation according to the official train-val split provided by the original ARKitScenes [4] dataset. For every scene, we created 3D point cloud associated per-point semantic labels in the original LabelMaker wordnet label space (186 classes). In some experiments, we project the dataset from wordnet label space to ScanNet200 label space to train it together with ScanNet200 dataset. We denote this converted dataset as ARKit LabelMakerSN200. To increase efficiency and make the experimental settings comparable to previous studies, we perform down-sampling on 3D meshes to a voxel size of 2 cm. Normal information is preserved and down-sampled simultaneously. Table 1 illustrates the scale of each dataset. ARKit LabelMaker dataset is the largest annotated real-world indoor semantic dataset.
Method Training Data Val mIoU Test mIoU
all head common tail all head common tail
MinkUNet [7]
vanilla ScanNet200 27.2 49.8 20.8 11.2 25.3 46.3 15.4 10.2
fine-tune (Ours) ARKit LabelMakerSN200 → ScanNet200 28.1 51.6 22.3 10.6 27.4 49.0 19.4 9.4
co-training (Ours) ARKit LabelMakerSN200 + ScanNet200 28.2 52.5 22.4 9.8 - - - -
PTv3 [40]
vanilla ScanNet200 35.2 56.5 30.1 19.3 37.8 - - -
fine-tune (Ours) ARKit LabelMakerSN200 → ScanNet200 36.4 56.7 31.2 21.6 - - - -
fine-tune (Ours) ARKit LabelMaker → ScanNet200 37.0 57.0 32.6 21.7 38.4 58.2 30.9 22.2
PPT [40] ScanNet200 + S3DIS + Structure3D → ScanNet200 36.0 - - - 39.3 59.2 33.0 21.6
PPT(Our ablation) ScanNet + ScanNet200 + Structure3D → ScanNet200 36.3 56.6 31.7 20.7 - - - -
PPT(Ours) ScanNet + ScanNet200 + ScanNet++ + Structure3D + ARKit LabelMaker 37.5 58.8 33.3 20.4 41.4 61.0 32.2 27.1
Table 2: 3D Semantic Segmentation Scores on ScanNet200 [30]. To investigate our large-scale dataset also helps with long-tail categories, we evaluate it on the ScanNet200 dataset. For both MinkowskiNet [7] and PointTransformerv3 [40], we compare it to vanilla training as well as training procedure proposed in [40]. We can show that common neural networks benefit from pre-training on automatically generated large-scale annotations. We also report the head, common and tail classes mean IoU.
@@
fine-tune (Ours) ARKit LabelMakerSN200 → ScanNet 74.1 -
@@
fine-tune (Ours) ARKit LabelMakerSN200 → ScanNet 78.9 -
fine-tune (Ours) ARKit LabelMaker → ScanNet 78.0 79.0
@@
PPT (Ours) ScanNet+ ScanNet200 + ScanNet++ + Structure3D + ARKit LabelMaker 79.1 79.8
@@
Table 3: 3D Semantic Segmentation Scores. Comparing training strategies for two top-performing models (PointTransformerV3 [40] and MinkowskiNet [7]) on ScanNet20 [9]. Adding ARKit LabelMakerSN200 through pre-training and co-training improves the performance for both models. With PonderV2 [49] and Mix3D [21], we compare large-scale pretraining to two other training strategies. Large-scale pre-training is superior to both, extensive data augmentation (Mix3D) and self-supervised pre-training (PonderV2).
Metrics.
We follow the standard metrics of the ScanNet 3D semantic segmentation task and compute the mean and per-class intersection-over-union (IoU), the mean per-class accuracy (mAcc) and the total per-point accuracy (tAcc).
4.3 Experiment Settings
Pre-training.
To investigate whether automatic labels are useful to learn strong features from imperfect annotations, we pre-train both, MinkowskiNet and PointTransformerV3, on our generated ARKit LabelMakerSN200 dataset. Afterwards, we fine-tune the pretrained models on the ScanNet and ScanNet200 dataset, respectively.
For PTv3 [40], we follow [40] employing the AdamW optimizer with OneCycleLR for 800 epochs of training. Similar to the fine-tuning of MinkowskiNet, we initially freeze the backbone and then solely train the classification head until convergence. Then, we fine-tune on ScanNet or ScanNet200 with a reduced learning rate of 0.0006. Besides ARKit LabelMakerSN200, we also pre-train PTv3 with ARKit LabelMaker in wordnet label space as the mapping from wordnet to ScanNet200 may reduce class diversity.
Co-training with ScanNet200.
With this experiment, we investigate whether ARKit LabelMaker can be seamlessly combined with existing datasets to increase dataset size and improve model performance. To this end, we merge ARKit LabelMakerSN200 with ScanNet200 and train a MinkowskiNet from scratch. Due to resource constraints, we conduct this experiment only with MinkowskiNet. The training setup follows the exact pre-training procedure described earlier for MinkowskiNet.
Joint-training.
We employ PTv3+PPT for joint training across multiple datasets and label spaces. In addition to ScanNet/ScanNet200, ScanNet++, and Structured3D, we incorporate our ARKit LabelMaker dataset. To maximize semantic class exposure, we adopt LabelMaker’s WordNet label space. Our training setup follows the exact PTv3+PPT configuration from [40], using the AdamW optimizer with a OneCycleLR scheduler and a learning rate of 0.05. We also integrate the LabelMaker WordNet label space into the normalization layer and final classification head.
4.4 Experiments
Pretrain Dataset fine-tune val mIoU
scratch 32.76
10% ARKit LabelMaker 33.30 (+0.54)
20% ARKit LabelMaker 35.29 (+2.53)
50% ARKit LabelMaker 36.29 (+3.53)
100% ARKit LabelMaker 37.04 (+4.28)
Structured3D (re-sampled to ARKit LabelMaker size) 32.21
Table 4: PTv3 [40] pretrained on different portion of ARKit LabelMaker dataset and Structured3D resampled to be the same size as ARKit LabelMaker. We then fine-tune on ScanNet200 and report validation mIoU. These experiments show the scaling potential of PTv3 and the advantage of real-world over synthetic data.
Method mIoU mAcc tAcc
ScanNet’s Annotation 17.7 21.3 70.6
LabelMakerV1 16.3 20.3 75.0
LabelMakerV2 w/o GSAM 17.6 21.0 77.1
LabelMakerV2 w/ GSAM 18.5 22.7 76.9
Table 5: Evaluation of LabelMaker on 5 ScanNet scenes. We use the same ScanNet scenes in the original LabelMaker’s paper with manual annotation in WordNet label space. Our LabelMakerV2 improves over ScanNet’s manual annotation.
In Table 3, we present the results for the ScanNet dataset. For MinkowskiNet [7], we show that pre-training on our large-scale, real-world ARKit LabelMakerSN200 dataset not only significantly improves the mean intersection-over-union compared to vanilla training but also outperforms other pre-training variants. Similar trends are observed in Table 2, which reports results on the ScanNet200 dataset. Co-training MinkowskiNet on ScanNet200 further confirms that our ARKit LabelMakerSN200 dataset enhances training without introducing a domain gap relative to ScanNet200.
Comparison to Self-supervision.
Table 3 shows that pre-training on our imperfect yet automatically generated labels outperforms self-supervised pre-training (PonderV2 [49]) and extensive data augmentation (Mix3D [21]). This highlights the importance of direct supervision with large-scale training data for effective 3D segmentation.
Comparison to Training on Synthetic Data.
Table 2 shows that PTv3 pre-trained on ARKit LabelMaker achieves comparable or superior performance to large-scale multi-dataset joint training. The approach in [40] relies heavily on Structure3D, a synthetically generated dataset, for pretraining, motivating a deeper analysis in Table 4. We compare pretraining on different subset sizes of ARKit LabelMaker to an equivalent amount of synthetic data from Structure3D [48]. Even at smaller scales, pretraining with real-world ARKit LabelMaker data proves more effective than synthetic data, which can even degrade performance at limited sizes. These findings strongly indicate that auto-labeling real-world recordings is significantly more effective per data point than relying on synthetic data.
Effect on Long-tail Classes.
For Point Transformer, integrating our ARKit LabelMaker into the joint training of PTv3+PPT results in a noticeable boost in validation and test mIoU. This version of PTv3 achieves state-of-the-art performance on both ScanNet and ScanNet200, with the latter benefiting significantly from improved tail class mIoU (Table 2, Figure 4). When comparing fine-tuned and vanilla models on the ScanNet200 validation set, our trained model shows a performance gain of +0.5% on head classes and +2.4% on tail classes. Compared to PTv3-PPT trained without ARKit LabelMaker, our model achieves a +0.8% gain on head classes and +5.5% on tail classes on the ScanNet200 test set. This effect persists even in a zero-shot setting on the Matterport3D dataset (Table 7). Including ARKit LabelMaker in training yields similar performance gains across Top-40 to Top-160 class sets. For MinkowskiNet, we do not observe a consistent improvement in tail class performance. On ScanNet200, validation mIoU slightly declines. In zero-shot evaluation on Matterport3D, the performance gain from adding ARKit LabelMakerSN200 over MinkowskiNet trained solely on ScanNet200 diminishes from Top-40 to Top-160 class sets.
LabelMakerV2 Method Variant mIoU mAcc tAcc
w/o GSAM w/o $g$-alignment 7.1 10.0 45.1
w/o GSAM   w/ $g$-alignment 10.3 13.4 70.6
w/ GSAM w/o $g$-alignment 7.9 11.0 44.9
w/ GSAM   w/ $g$-alignment 12.9 15.7 73.6
Table 6: We annotate three ARKitScenes scenes and perform ablation study of GSAM and gravity alignment. Our results shows that both GSAM and gravity alignment are helpful to the performance. This demonstrate the contribution of our novelty in improving LabelMaker’s pipeline. We also provide the visualization of ground truth labels and predictions of these scenes in Figure 5.
Figure 4: Correctly predicted tail class points on ScanNet200 validation set. We compare the number of correctly predicted points of selected tail class in ScanNet200 validation sets between PTv3 trained from scratch and the PTv3-PPT trained with our datasets. With our dataset, Point Transformer gains more ability to detect rase classes. Tail classes that are not predicted by any models are ignored in this plot, and we present the full tail class performance difference in the supplementary.
4.5 Ablation Studies
Different Training Regimes.
Between pre-training with ARKit LabelMaker and co-training with aligned labels, we observe no significant difference in Table 2. Co-training is primarily relevant for MinkowskiNet, as PTv3 allows for joint training across different output label spaces through PPT [40]. Increasing the dataset scale further through joint training on multiple datasets generally leads to better performance than pre-training solely on ARKit LabelMaker.
Evaluation of LabelMakerV2.
In Table 5, we evaluate our updated LabelMakerV2 pipeline on five ScanNet scenes with five manual annotations in LabelMaker’s WordNet label space, sourced from the original LabelMaker. The current pipeline achieves higher accuracy than even ScanNet’s original annotations. In Table 6, we conduct an ablation study to assess the performance improvements from integrating Grounded-SAM and gravity alignment. For this study, we manually annotate three ARKitScenes and provide additional visualizations in Figure 5. Figure 3 qualitatively demonstrates how gravity alignment corrects prediction errors in a video with continuous viewpoint rotation.
3D Scene
Manual Annotations
LabelMaker auto-labels (Ours)
OpenScene [23]
Figure 5: Visualization on ARKitScenes. From left to right: 3D scene, ground truth annotation (black regions indicate unannotated areas), LabelMaker annotations, OpenScene predictions.
Figure 6: Self-captured scenes and the semantic segmentation generated by LabelMakerv2. This figure provides a qualitative impression of the segmentation quality obtained by LabelMakerV2. The color mapping is defined here.
label space Top-40 NYU Classes Top-80 NYU Classes Top-160 NYU Classes
method mIoU mAcc tAcc mIoU mAcc tAcc mIoU mAcc tAcc
MinkowskiNet [7] trained on Matterport3D (supervised) - 50.8 - - 33.4 - - 18.4 -
MinkowskiNet trained on ScanNet200 33.1 43.5 74.5 19.9 28.4 72.3 11.7 17.5 71.5
MinkowskiNet trained on ARKit LabelMakerSN200 + ScanNet200 38.9 49.6 77.5 23.2 32.7 75.3 13.6 20.6 74.5
OpenScene[23] - 50.9 - - 34.6 - - 23.1 -
vanilla PTv3 trained on ScanNet200 36.2 46.1 77.3 22.0 29.8 75.0 13.0 18.8 74.1
PTv3-PPT without ARKit LabelMaker (Our reproduction) 41.4 51.2 79.9 27.0 36.2 77.7 16.3 24.4 76.8
PTv3-PPT trained with ARKit LabelMaker (Ours) 43.8 53.6 80.6 29.1 38.2 78.5 17.3 26.0 77.6
Table 7: Zero-shot evaluation on Matterport3D [6] test set region segmentation. We evaluate our trained PTv3 model on the ScanNet200 label space and map it to the top-40/80/160 NYU classes. Our models outperform fully-supervised MinkowskiNet and OpenScene.
ScanNet [9] ScanNet200 [30]
Method mAP25 mAP50 mAP mAP25 mAP50 mAP
vanilla PTv3 PointGroup (from [40] 77.5 61.7 40.9 40.1 33.2 23.1
vanilla PTv3 PointGroup (reproduced) 77.1 62.9 41.1 37.9 30.6 21.0
PTv3-PPT PointGroup pretrained [40] 78.9 63.5 42.1 40.8 34.1 24.0
PTv3-PPT PointGroup pretrained w/ ARKit LabelMaker (Ours) 78.5 62.6 41.4 42.9 34.9 24.5
Table 8: Instance segmentation scores on ScanNet(200) [30, 9]. Our dataset improves instance segmentation as a downstream task. Pre-training with our dataset increases performance on both ScanNet and ScanNet200, with a significant gain on ScanNet200.
Scaling Potential of LabelMaker and PTv3.
In Table 4, we conduct an ablation study by varying the number of pre-training samples from the ARKit LabelMaker dataset and evaluating fine-tuning performance on ScanNet200. The log-linear relationship between loss and data size is well-documented in large language models. In our experiments we cannot find such trend for validation loss at either the pre-training or fine-tuning stage. Instead, we find a log-linear relation between validation mIoU and dataset size from 20% onwards, which in turn indicates further scaling opportunities for even larger datasets ( Figure E3 ). We also run a pre-training experiment with the same data volume from the synthetic Structured3D dataset, which showed no performance gain. This further highlights the importance of our pipeline.
4.6 Transferability to other Domains
Efficacy in Downstream task.
Since ARKit LabelMaker is auto-labeled for 3D semantic segmentation, an important question is how useful this data generation approach is for other related 3D perception tasks. To assess this, we evaluate the general features learned through supervised training on our dataset by using PTv3 as the backbone and fine-tuning PointGroup [12] for 3D instance segmentation on ScanNet and ScanNet200. Results in Table 8 show a substantial performance improvement on ScanNet200 when incorporating our ARKit LabelMaker dataset. This suggests that auto-labeling as a method for scaling data is not limited to a single task but can serve as a general pretraining objective for learning robust features, akin to how ImageNet pretraining has become a standard for image encoders.
Zero-shot Evaluation on Matterport3D.
We perform inference on the Matterport3D [6] dataset using the ScanNet200 label space, and map the labels to the Top-40/80/160 NYU classes as done in [23]. PTv3 trained with ARKit LabelMaker surpasses both directly supervised models and OpenScene [23], an open-vocabulary model. These results indicate that ARKit LabelMaker has sufficient scale and diversity of real-world appearance to strongly improve model generalization.
Data Scaling via Mobile Integration.
While LabelMaker demonstrates success in large-scale 3D training on ARKitScenes, the dataset’s diversity remains limited. To scale our pipeline to arbitrary real-world environments, we integrate the iOS app Scanner 3D into LabelMaker, leveraging modern mobile devices’ ubiquitous RGB-D capture capabilities. This integration enables automatic annotation of scenes recorded with consumer iPhones. We validate scalability by processing two self-captured scenes, a kitchen and a fireplace, recorded using an iPhone 12 Max in a holiday cottage. Figure 6 showcases reconstructed scenes and their semantic segmentations, confirming the pipeline’s effectiveness for diverse real-world settings.
4.7 Limitations & Broader Impact
While we enhance LabelMaker [37] with an improved point cloud pipeline, we omit the generation of 2D segmentation maps. The computational cost of NeRF-based lifting across the entire ARKitScenes dataset exceeds our available resources, requiring approximately 12 additional GPU hours per scene. An interesting future research direction would be to explore more efficient 2D lifting methods and assess whether training 2D models on this data yields performance gains comparable to those observed for 3D models. Furthermore, 20 ARKitScenes scenes are excluded from processing due to missing pose data. Since LabelMakerV2 relies on accurate poses, future iterations of the software stack could incorporate techniques to reconstruct missing pose information, enabling broader dataset coverage. Like the original LabelMaker [37], our improved pipeline does not achieve perfect accuracy. While [37] demonstrated that its annotations are comparable to crowd-sourced human labels, training on noisy data always carries the risk of introducing systematic errors. For safety-critical applications, rigorous evaluation on accurately annotated data remains essential when leveraging tools like ours for training data generation. Does large-scale pretraining with automatic labels exhibit similar trends as seen in language and image generation tasks? Our results suggest so, showing measurable improvements across multiple architectures and tasks when (pre)training on ARKit LabelMaker. However, while real-world data proves significantly more effective than synthetic data, the scale of generated 3D data remains orders of magnitude smaller than that of image datasets. Our pipeline facilitates data collection, making it easier to expand training datasets as more scans become available.
5 Conclusion
In this paper, we introduced ARKit LabelMaker, the largest real-world 3D RGB-D dataset with 3D semantic annotations, generated automatically using an enhanced version of LabelMaker [37], which we call LabelMakerV2. While these labels are inherently imperfect due to automation, we demonstrate their effectiveness in pre-training widely used 3D semantic segmentation models, significantly outperforming traditional, self-supervised, and augmentation-heavy training approaches. This aligns with trends in language and image generation, where scaling up training data has led to substantial performance gains. To support further data collection for training and evaluation, we also provide integration with an existing iOS app.
Acknowledgements.
We thank Xiaoyang Wu, author of PointTransformerV3 [40], for valuable and helpful advice. This project is partially supported by an SNSF Postdoc.Mobility fellowship, the ETH Foundation through a CareerSeed grant, and the Lamarr Institute.
Supplementary Material
A Dataset Class Statistics
Dataset Statistics of ARKit LabelMaker. In Figure A1, we present the point count for each class in the LabelMaker WordNet label space. Our dataset maintains a substantial data distribution even across tail classes.
Figure A1: Number of points for each ARKitLabelMaker class.
B PTv3 Results on ScanNet++
We also report the training and evaluation results of PTv3 on ScanNet++ in Table B1. Unfortunately, the numbers are not fully comparable, because we were so far unable to reporduce the validation results of PTv3. When the authors of [40] released PTv3’s performance on ScanNet++, they expanded Structured3D’s training set from 6,519 to 18,348 samples, which we refer to as Structured3D v2. Due to limited computational resources, we could not train with this updated version of Structured3D yet. We will update ScanNet++ results once the new result is available. Our pre-training and joint-training (PPT) experiments show performance gains over vanilla PTv3, with PTv3-PPT achieving similar improvements to the original PTv3-PPT but with significantly less training data.
PTv3 Variant Training Data #Data val mIoU
vanilla ScanNet++ 713 41.8
fine-tune (Ours) ARKit LabelMakerSN200 → ScanNet++ 4471 → 713 42.5
PPT [40] ScanNet200 + ScanNet++ + Structure3Dv2 45868 45.3
PPT (Ours) ScanNet200 + ScanNet++ + ARKit LabelMaker 11168 44.5
PPT (Ours) ScanNet+ ScanNet200 + ScanNet++ + Structure3D + ARKit LabelMaker 30386 44.6